40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 12 03:52:51 2011, kiwamu, Update, SUS, PRM OSEM adjustment 
    Reply  Sat Aug 13 01:54:18 2011, kiwamu, Update, SUS, PRM OSEM adjustment part II before.pngafter.png
Message ID: 5203     Entry time: Fri Aug 12 03:52:51 2011     Reply to this: 5219
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: PRM OSEM adjustment 

[Suresh / Kiwamu]

 We tried adjusting the OSEMs on PRM, but we didn't complete it due to a malfunction on the coils.

The UL and LL coils are not working correctly, the forces are weak.

Tomorrow we will look into the satellite box, which is one of the suspects.

 

 During the adjustment we found that the POS excitation force was unequal in each sensor.

At the beginning we thought it's because of the difference of the sensitivity in each OSEM due to the bad OSEM orientations.

However it turned out that it comes from the actual force imbalance on each coil.

We checked the force of each coil by putting an offset (-2000 cnts) in each output digital filter and looked at the OSEM signals in time series.

The UL and LL coils are too weak and the responses are almost buried in the noise of the OSEMs in time series.

We briefly checked some analog electronics and found the DAC, AI board and deWhitening board were healthy.

We were able to see the right amount of voltage from the monitor pin on the front panel of the coil driver.

So something downstream are suspicious, including the satellite box, feedthrough and coils.

- - -

Although the coil issue, it could be worth trying to check the input matrix.

ELOG V3.1.3-